OpenStudy (casodi):

Differentiate with respect to X sin2x/2cosx

OpenStudy (dayakar):

sin2x/2cosx \[=\frac{ 2 \sin x cosx }{ 2 \cos x }\] [ since sin 2x = 2sinxcosx ]

OpenStudy (dayakar):

d/dx(sinx) = ? [ 2 cos x cancelled]

OpenStudy (dayakar):

do u understand

OpenStudy (casodi):

Oh i understand now. the sinx is left and that gets derived

OpenStudy (dayakar):

:)